OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8.

OpenStudy (luffingsails):

y varies directly as the square of x means: y = kx^2 (where we aren't sure what k is yet) You are given y = k*3^2 = k*9 = 54 So, k = 6. Now let x = 8 y = 6 * 8^2 = 6 * 64 = 384.

OpenStudy (luffingsails):

Make sense?

OpenStudy (anonymous):

yup thank you:)

OpenStudy (luffingsails):

Cool.